# Multimodal AI predicts clinical outcomes of drug combinations from preclinical data

### v1

@@ Abstract
Predicting clinical outcomes from preclinical data is essential for identifying safe and effective drug combinations, reducing late-stage clinical failures, and accelerating the development of precision therapies. Current models rely on structural or target-based features to identify high-efficacy, low-toxicity drug combinations. However, these approaches fail to incorporate the multimodal data necessary for accurate, clinically-relevant predictions. Here, we introduce Madrigal, a multimodal AI model that learns from structural, pathway, cell viability, and transcriptomic data to predict drug combination effects across 953 clinical outcomes and 21,842 compounds, including combinations of approved drugs and novel compounds in development. Madrigal uses a transformer bottleneck module to unify preclinical drug data modalities while handling missing data during training and inference—a major challenge in multimodal learning. It outperforms single-modality methods and state-of-the-art models in predicting adverse drug interactions. Madrigal performs virtual screening of anticancer drug combinations and supports polypharmacy management for type II diabetes and metabolic dysfunction-associated steatohepatitis (MASH). It identifies transporter-mediated drug interactions. Madrigal predicts resmetirom, the first and only FDA-approved drug for MASH, among therapies with the most favorable safety profile. It supports personalized cancer therapy by integrating genomic profiles from cancer patients. Using primary acute myeloid leukemia samples and patient-derived xenograft models, it predicts the efficacy of personalized drug combinations. Integrating Madrigal with a large language model allows users to describe clinical outcomes in natural language, improving safety assessment by identifying potential adverse interactions and toxicity risks. Madrigal provides a multimodal approach for designing combination therapies with improved predictive accuracy and clinical relevance.
@@
Combination therapies improve clinical outcomes by enhancing efficacy and reducing adverse effects through lower doses or complementary mechanisms [1, 2, 3]. While drug combinations can enhance treatment efficacy, they also elevate the risk of adverse drug reactions due to drug-drug interactions (DDIs). For example, the combination of infliximab and azathioprine has been shown to be more effective than monotherapy in inducing and maintaining corticosteroid-free remission in Crohn’s disease, but has also been shown to be associated with infections and other adverse effects [4]. Similarly, combining nivolumab with chemotherapy has improved overall survival in advanced esophageal squamous cell carcinoma, yet this approach is associated with higher incidences of treatment-related adverse events [5]. These interactions can pose significant risks for vulnerable populations such as cancer survivors [6], patients with chronic [7] and neurological diseases [8]. Despite the potential of combination therapies, identifying clinically effective and safe drug combinations remains challenging due to the overwhelming number of possible drug pairings and the heterogeneity of clinical effects [9, 10].
@@
A major challenge in predicting drug combination effects for novel compounds—those in preclinical or early clinical development—is the absence of critical data that emerge only in later stages. These missing data, including clinical safety profiles, long-term efficacy, and pharmacokinetics, constrain accurate prediction of drug interactions and therapeutic effects. Accurately predicting drug combinations with limited preclinical data is thus essential for improving clinical success rates [11], reducing patient risk and prevent unnecessary clinical investigations.
@@
Various preclinical data modalities provide insight into the clinical effects of drug combinations, including molecular structures, mechanisms of action, and perturbation outcomes from cell-based assays. Although molecular structure is universally available and provides insights into drug bioactivity [12, 13], it is often insufficient for characterizing drug combinations. This limitation arises from complex pharmacodynamic interactions that extend beyond molecular properties [14] (Supplementary Fig. S1a) and requires a broader understanding of drug action[15]. Perturbation-based modalities include transcriptional changes in cell lines [16] and cell viability profiles [17] following chemical perturbations, which can be measured at high throughput [16] and capture complex biological responses to drugs [18, 19, 20]. Despite their relevance, perturbation data remain underutilized in predictive modeling, yet they are essential for understanding drug synergy and safety [21, 22]. Transcriptional phenotypes provide readouts of drug-protein interactions and pathway-level changes in cellular activity [23, 24, 25, 26]. Cell viability data reveal how drugs influence signaling pathways across tissues and cell types, enabling the identification of gene functions relevant to drug action [27, 28, 29]. This is critical for target prioritization[30], detecting adverse drug interactions[31, 32, 33], and managing polypharmacy in patients with comorbidities.
@@
A key barrier to utilizing preclinical data to achieve accurate prediction of drug combination clinical outcomes is the “missing modality” problem[34, 35, 36] (Fig. 1c), where crucial data are unavailable for novel compounds in preclinical stages and even for some approved drugs. Approaches that assume complete data availability during training cannot leverage drugs with incomplete modality profiles. This disproportionately affects novel compounds lacking pathway annotations and early-stage experimental drugs with sparse toxicity data. To address this, there is a need for multimodal AI models that can handle missing modalities at both training and inference stages of AI modeling.
Here, we introduce Madrigal, a multimodal AI model that predicts drug combination outcomes. Madrigal unifies drug structural, pathway, cell viability, and transcriptomic modalities throughout the AI model, including combinations of compounds that lack preclinical data modalities or reported clinical efficacy and safety data (Fig. 1c). By aligning modality-specific drug representations using contrastive learning [37, 38], Madrigal learns a unified multimodal latent space for improved performance across clinical outcomes and can be integrated with large language models (LLMs) to model outcomes beyond those included in predefined medical vocabularies and standardized ontologies. We use Madrigal to evaluate new and approved cancer combination therapies involving poly (ADP-ribose) polymerase inhibitors (PARPi), including olaparib, talazoparib, and rucaparib. Madrigal predicts olaparib + abiraterone to have a more favorable safety profile than olaparib + paclitaxel, based on lower predicted adverse effect score (a “favorable safety profile” is defined based on predicted score of adverse effects). Beyond cancer treatments, we also evaluated Madrigal predictions for treating chronic metabolic disorders, resulting in insights into complex drug regimens used in type II diabetes and metabolic dysfunction-associated steatohepatitis (MASH). Madrigal supports personalized cancer therapy by predicting effective drug combinations using patient genomic profiles from primary acute myeloid leukemia samples [39, 40] and patient-derived tumor xenografts [41]. It identifies responders by modeling drug synergy and treatment response across diverse genetic backgrounds.
@@ Results
Predicting effective combination therapies requires models that can generalize across diverse data modalities available. Existing approaches often assume complete modality data during both training and inference, limiting their applicability when certain modalities are missing—a common challenge in preclinical and clinical studies. This constraint is particularly problematic for predicting drug combination outcomes, as information on novel or experimental compounds is often sparse. Madrigal is a universal multimodal model designed to handle incomplete modality inputs during both training and inference. It predicts clinical outcomes and adverse reactions of drug combinations using preclinical data, supporting decision-making for both existing combination therapies and novel drug candidates. Using structural drug information, molecular pathways knowledge graph, transcriptomic responses, and cell viability data, it predicts combination therapy effects across 953 outcomes (Fig. 1a,b, such as “increase in QTc prolongation”) and through integration with an LLM, Madrigal-LLM extends to free text descriptions of outcomes beyond codified clinical vocabularies (Fig. 6a).
For each pair of drugs, Madrigal processes available modality inputs individually via modality-specific encoders, mapping each modality to a unified embedding space. These modality-specific embeddings function as tokens (vector representations of input data) that are fed into a fusion module comprising multiple customized transformer encoder layers, concluding with a cross-attention layer (Supplementary Fig. S1b). Bottleneck tokens are placed between structure, pathway, cell viability, and transcriptomic modality tokens, balancing information from the distinct transcriptomic responses of each cell line[42]. The fusion module generates a single embedding by applying cross-attention between a summarization query token and the bottleneck tokens from the last transformer encoder block [43, 44]. The resultant multimodal embedding for each drug, when paired and passed through a decoder module, outputs a final predicted score for each outcome (Fig. 1b, Supplementary Fig. S1b).
To ensure alignment between data modalities, Madrigal uses contrastive pretraining to align modality-specific embeddings with the structure modality, the only universal data type available for all small molecule compounds (Fig. 1a,c). After alignment, Madrigal generates unified representations in latent space, allowing for fine-tuning of drug combination datasets (formatted as [drug 1, drug 2, outcome]; Methods Sec. 2) (Fig. 1d, Supplementary Fig. S1a).
Evaluating Madrigal across drug combination prediction tasks.
We first evaluated Madrigal on tasks in which (1) all samples involving a specific drug combination (“split-by-drug pairs”) or (2) all samples involving a specific drug (along with its combinations) were held out during training (“split-by-drugs”). The latter scenario better simulates the cases where the outcomes need to be predicted for a novel compound combined with an approved drug, as the drug that is held out mimics new compounds not encountered by the model during training (Fig. 2a).
Madrigal model is trained on the following two datasets: TWOSIDES (2019-11-15), a self-reported DDI dataset extracted from FDA Adverse Event Reporting System (FAERS) [45] containing 4,656,138 drug combinations across 1,457 drugs and 795 outcomes, and DrugBank (2023-01-04), an expert-curated dataset with 1,188,371 drug combinations involving 3,632 drugs and 158 outcomes [46] (Methods Sec. 1). To test Madrigal’s generalizability, we created several dataset splits: in addition to random splits by drug pairs or drugs, we designed two splits to test model generalization. In the split-by-drugs (target) approach, the drugs in the test set have none or only a few therapeutic targets in common with those in the training set. In the split-by-drugs (ATC) approach, drugs from specific first level categories within the Anatomical Therapeutic Chemical (ATC) classification system were excluded from the training set (Methods Sec. 1). These splits significantly increased structural differences between training and test drugs compared to random splits (Supplementary Fig. S2a). For each dataset and splitting strategy, models are trained separately on the respective dataset and split for evaluation. Unless otherwise specified, the model trained on the full DrugBank dataset is used for pharmacological applications throughout this paper.
These dataset splits reflect challenges in drug combination development, where novel compounds are developed for new targets, often with limited preclinical data. While models have performed well on typical drug combination benchmarks [47, 46, 45], they usually assume all possible drugs have been encountered during training, an assumption that does not match scenarios in which new compounds are paired with existing drugs. For example, an approved drug may be combined with a novel compound to mitigate safety issues, enhance efficacy, or expand the approved drug’s indications. The ability to model interactions with such novel compounds early in development, when only limited preclinical data is available, could streamline the development of combination therapies [11].
In our comparisons, we included state-of-the-art models categorized by data modalities: chemical structure-based models (DeepDDI [48], CASTER [49], GMPNN-CS [50]) that capture molecular structure through learning on molecular graphs; knowledge graph (KG)-based models (DDKG [51]) that capture drug mechanism of action and include information on targets, enzymes, transporters, drug indications, and adverse effects; and multimodal models (MUFFIN [52], TIGER [53]). These models use either late fusion [51, 52, 53]—where each drug molecule is encoded separately and merged—or early fusion, where molecular interactions are modeled from the start [48, 49, 50]. For evaluation metrics, we used the area under the receiver-operating curve (AUROC), the area under the precision-recall curve (AUPRC), and the maximum F measure (Fmax) (Methods Sec. 3.1). To further reflect drug combination development when information is limited for novel compounds, we restricted Madrigal’s input during testing to only those modalities typically available preclinically (Fig. 2a). In contrast, baselines were provided with full multimodal inputs. This deliberate limitation makes the prediction task more challenging for Madrigal.
Madrigal predicts multimodal drug combination safety.
Madrigal demonstrated strong performance across both datasets, various types of splitting, and evaluation metrics (split-by-drugs (target) setting in Fig. 2b, split-by-drugs (ATC, random) and split-by-drug pairs settings in Supplementary Fig. S3). While GMPNN-CS matched Madrigal’s AUROC (+0.001) in DrugBank (split-by-drugs target) and CASTER excelled in split-by-drug pairs, Madrigal was the most reliable overall, particularly in challenging scenarios.
In the split-by-drugs (target) setting, Madrigal demonstrated an average improvement of 10.7% across all metrics compared to structure-based models on TWOSIDES. Similarly, on DrugBank, Madrigal outperformed most structure-based models across all metrics by, on average, 6.2% (Fig. 2b). Although CASTER and GMPNN-CS utilize explicit substructure learning strategies and employ early fusion to enable information flow between drugs early on, Madrigal adopts a late fusion strategy, first integrating multiple modalities for each compound before combining the drug representations, all while using a simple lightweight structural encoder. The superior performance of Madrigal suggests that its effectiveness is likely due to the efficient alignment of modalities rather than relying solely on modality-specific encoding. This approach allows Madrigal to leverage information from various data sources more effectively, improving predictive capabilities for drug combinations.
Madrigal improved upon multimodal (KG-structure) models by 22.5% and 12.8% in AUROC in the two datasets on average, under this challenging split (Fig. 2b). Similar performance gains are observed across other dataset splits, demonstrating that Madrigal advances drug combination outcome prediction through the integration and alignment of diverse modalities (Supplementary Fig. S3). Ablation studies validated the contributions of contrastive learning and multimodal learning across both datasets compared to using only structural information, showing that Madrigal’s performance benefits from including multiple data types (Fig. 2b, Supplementary Fig. S3).
To evaluate the robustness of Madrigal, we examined its predictive precision in drug safety datasets. Our analysis hypothesizes that test drugs with greater representation and structural similarity to drugs in the training set yield a higher prediction accuracy for DDIs. We observed that increased structural similarity enhances DDI prediction accuracy, an effect seen not only in the complete model (+modality alignment) but also in ablation models without alignment (+multimodal) and those relying on structure only (Fig. 2c). The similarity of the target profile further reinforces these predictions, highlighting the role of underlying biological mechanisms in drug safety modeling (Fig. 2d).
@@
The multimodal approach in Madrigal outperforms the unimodal model across diverse outcomes, particularly when the modalities are aligned (Fig. 2e). Predictions for narrowly defined clinical outcomes, which correspond closely to specific biological pathways, are generally more accurate than those for broader phenotypes, underscoring the importance of pathway-specific knowledge (Supplementary Fig. S2b) [47]. Performance improvements are notable as additional modalities are integrated (Supplementary Fig. S2c). Attention weight analysis reveals that cell viability data contribute minimally, while transcriptomics plays a significant role despite its lower prevalence (Supplementary Fig. S2d).
Madrigal predicts drug safety profiles and captures transporter-mediated interactions.
@@
To assess whether Madrigal can capture safety signals beyond drug combination contexts, we tested it on individual drugs by pairing each drug with itself as input. Madrigal predictions correlate with established safety profiles for specific outcomes, such as liver injury (Drug-Induced Liver Injury, DILIrank) [54], cardiotoxicity (Drug-Induced Cardiotoxicity, DICTrank) [55], and QT prolongation (Drug-Induced QT Prolongation Atlas, DIQTA) [56] (Fig. 3a-c, Supplementary Fig. S4). This shows the capability of Madrigal to identify clinical safety signals, even when applied to single drugs.

### v2

@@ Abstract
Predicting clinical outcomes from preclinical data is essential for identifying safe and effective drug combinations, reducing late-stage clinical failures, and accelerating the development of precision therapies. Current AI models rely on structural or target-based features but fail to incorporate the multimodal data necessary for accurate, clinically relevant predictions. Here, we introduce Madrigal, a multimodal AI model that learns from structural, pathway, cell‑viability, and transcriptomic data to predict drug‑combination effects across 953 clinical outcomes and 21,842 compounds, including combinations of approved drugs and novel compounds in development. Madrigal uses an attention bottleneck module to unify preclinical drug data modalities while handling missing data during training and inference, a major challenge in multimodal learning. It outperforms single‑modality methods and state‑of‑the‑art models in predicting adverse drug interactions, and ablations show both modality alignment and multimodality are necessary. It captures transporter-mediated interactions and aligns with head-to-head clinical trial differences for neutropenia, anemia, alopecia, and hypoglycemia. In type 2 diabetes and MASH, Madrigal supports polypharmacy decisions and prioritizes resmetirom among safer candidates. Extending to personalization, Madrigal improves patient-level adverse-event prediction in a longitudinal EHR cohort and an independent oncology cohort, and predicts ex vivo efficacy in primary acute myeloid leukemia samples and patient‑derived xenograft models. Madrigal links preclinical multimodal readouts to safety risks of drug combinations and offers a generalizable foundation for safer combination design.
@@
Combination therapies are central to modern treatment: by leveraging complementary mechanisms or dose reduction, they can enhance efficacy, mitigate single-agent toxicities, or both [1, 2, 3]. Yet combining agents also increases the risk of adverse drug reactions driven by drug-drug interactions (DDIs). For example, infliximab plus azathioprine improves corticosteroid-free remission in Crohn’s disease compared with monotherapy but is associated with infections and other adverse events [4]; likewise, nivolumab with chemotherapy prolongs overall survival in advanced esophageal squamous cell carcinoma at the cost of higher treatment-related toxicities [5]. These risks are magnified in vulnerable populations, including cancer survivors [6] and patients with chronic [7] and neurological diseases [8]. Moreover, identifying combinations that are both effective and safe remains challenging due to the combinatorial explosion of possible pairs and the heterogeneity of clinical effects across patients and indications [9, 10].
@@
A major challenge in predicting drug combination effects for novel compounds—those in preclinical or early clinical development—is the lack of critical data that emerge only in later stages of testing. Missing information such as clinical safety profiles, long-term efficacy, and pharmacokinetics limits the ability to accurately forecast drug interactions and therapeutic responses. Developing predictive models that can reliably infer combination outcomes from limited preclinical data is therefore essential for improving clinical success rates [11], minimizing patient risk, and avoiding unnecessary clinical studies.
@@
Various preclinical data modalities provide complementary insight into the clinical effects of drug combinations, including molecular structures, mechanisms of action, and perturbation outcomes from cell-based assays. Although molecular structure is universally available and informative for drug bioactivity [12, 13], it is often insufficient to fully characterize combinations. This limitation stems from complex pharmacodynamic interactions that extend beyond intrinsic molecular properties [14] (Supplementary Fig. S1c) and demands a broader understanding of drug action[15]. Perturbation-based modalities, such as transcriptional changes in cell lines [16] and cell-viability profiles [17] following chemical perturbations, can be measured at high throughput and capture complex biological responses to drugs [18, 19, 20]. Despite their relevance, perturbation data remain underutilized in predictive modeling, yet they are essential for understanding drug synergy and safety [21, 22]. Transcriptional phenotypes provide readouts of drug-protein interactions and pathway-level changes in cellular activity [23, 24, 25, 26], while cell-viability data reveal how drugs influence signaling pathways across tissues and cell types, enabling the identification of gene functions relevant to drug action [27, 28, 29]. This information is critical for target prioritization [30], detecting adverse drug interactions [31, 32, 33], and managing polypharmacy in patients with comorbidities.
@@
A key barrier to predicting clinical outcomes of drug combinations from preclinical data is the “missing-modality” problem [34, 35, 36] (Fig. 1c, Supplementary Fig. S1d), in which crucial data are unavailable for novel compounds in preclinical or early clinical stages and even for some approved drugs. Methods that assume complete data during training either discard or distort drugs with incomplete modality profiles, limiting generalization to real-world settings. This disproportionately affects novel compounds lacking pathway annotations and early-stage experimental drugs with sparse toxicity data. Accordingly, there is a need for multimodal AI models that are robust to missing modalities at both training and inference, leveraging available information without relying on complete profiles.
Here, we introduce Madrigal, a multimodal AI model that predicts drug-combination outcomes (safety-related phenotypes). Madrigal integrates structural, pathway, cell viability, and transcriptomic modalities for each drug and operates when some modalities are absent at both training and inference (Fig. 1c). Using contrastive learning[37, 38], it maps modality-specific drug encodings into a shared latent space that preserves pharmacologic signal and allows observed modalities to inform missing ones. For a candidate drug pair, Madrigal then fuses the aligned representations through an attention bottleneck to generate risk scores for clinical outcomes. Madrigal captures transporter-mediated interactions: combinations that share transporters show higher predicted risks for serum-level and excretion-related outcomes, and exemplar pairs (e.g., doxycycline with tacrolimus) rank among the most concerning in transporter-relevant phenotypes. We test Madrigal in predicting safety of combination therapies in head-to-head clinical trials. In 35 post-2000 advanced-stage trials with multiple combination arms, Madrigal’s ordering of predicted risk agrees with observed differences for key adverse events (alopecia, anemia, hypoglycemia, neutropenia), indicating alignment between predicted outcomes and prospectively measured toxicities. We further evaluate Madrigal for chronic metabolic disorders, including type 2 diabetes (T2D) and metabolic dysfunction-associated steatohepatitis (MASH), where it supports polypharmacy management and ranks resmetirom, the first FDA-approved drug for MASH, among candidates with the most favorable predicted safety profiles. Madrigal supports personalized therapies by predicting effective combinations using patient demographic and genomic profiles from primary acute myeloid leukemia samples [39, 40] and patient-derived tumor xenografts [41]. In patient datasets, Madrigal predicts patient-level adverse-event risk of drug combinations in a longitudinal event-time cohort [42] and an independent oncology cohort at a major cancer center.
@@ Results
Predicting effective combination therapies requires models that generalize across diverse data modalities and remain reliable when some modalities are unavailable. Many existing approaches assume complete modality coverage during training and inference, which limits their utility in preclinical and clinical settings where information on novel or experimental compounds is often sparse.
@@
Madrigal is a multimodal model designed to handle incomplete modality inputs during both training and inference. It predicts clinical outcomes and adverse reactions of drug combinations from preclinical data, supporting decisions for both existing therapies and new candidates. Using structural drug information, a molecular pathways knowledge graph, transcriptomic responses, and cell viability data, Madrigal predicts combination effects across 953 outcomes (for example, “increase in QTc prolongation”; Fig. 1a; Supplementary Note S6).
For each drug pair, Madrigal encodes each modality with a modality-specific encoder and maps the resulting embeddings into a unified space through an attention bottleneck fusion module (Supplementary Fig. S1b, Supplementary Note S3). Bottleneck tokens are inserted between structure, pathway, cell viability, and transcriptomic embeddings to regulate information flow and to balance signal from transcriptomic responses across cell lines [43]. The fusion module produces a single multimodal embedding via cross-attention between a summarization query token and the bottleneck tokens from the last attention bottleneck layer [44, 45]. Pairwise combination of the two drug embeddings followed by a prediction head yields a score for each outcome (Fig. 1b; Supplementary Fig. S1a). To align data modalities, Madrigal uses contrastive learning that anchors modality-specific embeddings to the structure modality, which is universally available for small molecules (Fig.1a,c). After alignment, Madrigal forms unified latent representations that enable fine-tuning on drug combination datasets (Methods Sec. 2; Fig. 1d; Supplementary Fig. S1a). This design preserves predictive performance when some modalities are missing at inference and improves training efficiency.
Benchmarking Madrigal across challenging drug-combination tasks.
We evaluate Madrigal on two settings: (1) holding out all samples for specific drug pairs (“split-by-drug pairs”) and (2) holding out all samples for specific drugs together with any of their pairings (“split-by-drugs”). The split-by-drugs setting better reflects prediction for a novel compound combined with an approved partner (Fig. 2a).
Madrigal is trained on two datasets: TWOSIDES (2019-11-15), a FAERS-derived resource with 4,656,138 combinations across 1,457 drugs and 795 outcomes [46], and DrugBank (2023-01-04), an expert-curated resource with 1,188,371 combinations involving 3,632 drugs and 158 outcomes [47] (Methods Sec. 1). To probe Madrigal’s generalization, we introduce two harder variants of split-by-drugs. In split-by-drugs (target), test-set drugs share minimal therapeutic targets with training drugs. In split-by-drugs (ATC), we exclude drugs of certain first-level Anatomical Therapeutic Chemical (ATC) categories from training (Methods Sec. 3.1). Both strategies increase structural dissimilarity between training and test drugs relative to random splits (Supplementary Fig. S2a). For each dataset and split, models are trained separately and evaluated on the corresponding test set.
These splits mirror real development scenarios in which a novel compound is combined with an approved drug to mitigate safety issues, enhance efficacy, or extend indications despite limited preclinical data for the new agent. We compare against state-of-the-art models spanning three modality classes: structure-based models (DeepDDI [48], CASTER [49], GMPNN-CS [50]); knowledge-graph models (DDKG [51]); and multimodal models (MUFFIN [52], TIGER [53]). We report area under receiver-operator curve (AUROC), area under precision-recall curve (AUPRC), and maximum F measure (Fmax) (Methods Sec. 3.3).
To reflect real constraints when information is sparse for novel compounds, we restrict Madrigal’s test-time inputs to modalities typically available preclinically (Fig. 2a). In contrast, other models receive their full multimodal inputs. This deliberate asymmetry makes the task harder for Madrigal and directly assesses robustness to missing modalities while preserving a fair, task-matched comparison.
Performance across challenging, clinically realistic settings.
[…]
Relative to multimodal KG-structure models, Madrigal improves AUROC by 22.5% and 12.8% on average on TWOSIDES and DrugBank, respectively, under split-by-drugs (target) (Fig. 2b). Similar gains hold across split-by-drugs (ATC), split-by-drugs (random), and split-by-drug pairs, indicating that integrating and aligning diverse modalities advances drug-combination outcome prediction (Supplementary Fig.S3). Ablations confirm that both contrastive modality alignment and multimodality contribute beyond structure alone (Fig.2b; Supplementary Fig. S3).
We next examine robustness. We hypothesize that test drugs with greater representation and higher structural similarity to training drugs will yield higher accuracy. Accuracy increases with structural similarity for the full model, for the model without contrastive alignment (w/o CL), and for the structure-only ablation (Fig. 2c). Similarity in target profiles further strengthens performance (Fig. 2d).
@@
Multimodal Madrigal outperforms the unimodal model across outcome types, with the largest gains when modalities are aligned (Fig. 2e). Predictions for narrowly defined outcomes that map to specific biological pathways are generally more accurate than for broader phenotypes, consistent with the value of pathway-resolved knowledge (Supplementary Fig. S2b) [54]. Performance improves as additional modalities are incorporated (Supplementary Fig. S2c,e). Including an additional bioassay modality further improves performance across datasets and splits (Supplementary Table S9; Supplementary Note S8). Attention-weight analyses indicate that transcriptomic signals contribute strongly despite lower prevalence (Supplementary Fig. S2d).
Single-drug safety profiling and transporter-mediated interactions.
@@
To assess whether Madrigal captures safety signals beyond combination contexts, we evaluate it on individual drugs by pairing each drug with itself. Madrigal’s predictions correlate with established safety profiles for liver injury (DILIrank) [55], cardiotoxicity (DICTrank) [56], and QT prolongation (DIQTA) [57] (Fig. 3a-c; Supplementary Fig. S4), indicating that Madrigal surfaces clinically relevant single-drug risks from preclinical modalities.